find the momentum of an electron with a wavelength of 456 no
1
Expert's answer
2017-12-22T04:37:54-0500
Answer:
p = h/ λ
h = 6.62606957(29)×10^−34 J·s.
p = 6.62606957(29)×10^−34 / 456 ·10-9 = 0.014·10^-25